Solve:
6(x-2)-2(x-1)=2

I at first tried to use the distributive property... but I don't believe it turned out so well!
Help, please?
Thank you!


Answer by nerdybill(5411)

6(x-2)-2(x-1)=2
Distributing:
(6x-12)-(2x-2)=2
Distributing the "negative" sign to terms inside right parenthesis:
(6x-12)-2x+2=2
6x-12-2x+2=2
4x-10=2
4x=12
x = 3